Reports 1-1 of 1 Clear search Modify search
AdV-TCS (Ring Heater)
taranto - 0:52 Friday 01 July 2022 (56320) Print this report
NI RH switch on for CH centering measurement

After the ISC shift (Entry 56318), at 20.52 UTC the NI-CH was switched off in order to cool down the CP.

At 22.19 UTC HWS-DET acquisition started and at 22.27 UTC NI-RH  was switched ON.

Search Help